2.Yowell Company began operations on January 1, Year 1. During Year 1, the company engaged in the following cash transactions:

  1. 1) issued stock for $46,000
  2. 2) borrowed $28,000 from its bank
  3. 3) provided consulting services for $44,000 cash
  4. 4) paid back $18,000 of the bank loan
  5. 5) paid rent expense for $10,500
  6. 6) purchased equipment for $15,000 cash
  7. 7) paid $3,300 dividends to stockholders
  8. 8) paid employees' salaries of $24,000


What is Yowell's net cash flow from operating activities?

Multiple Choice

  • Inflow of $6,200

  • Inflow of $9,500

  • Inflow of $37,500

  • Inflow of $20,000

3.Yowell Company began operations on January 1, Year 1. During Year 1, the company engaged in the following cash transactions:

  1. 1) issued stock for $70,000
  2. 2) borrowed $40,000 from its bank
  3. 3) provided consulting services for $68,000 cash
  4. 4) paid back $30,000 of the bank loan
  5. 5) paid rent expense for $16,500
  6. 6) purchased equipment for $27,000 cash
  7. 7) paid $4,500 dividends to stockholders
  8. 8) paid employees' salaries of $36,000


What is Yowell's notes payable balance at the end of Year 1?

Multiple Choice

  • $40,000

  • $30,000

  • $0

  • $10,000

4.Yowell Company began operations on January 1, Year 1. During Year 1, the company engaged in the following cash transactions:

  1. 1) issued stock for $64,000
  2. 2) borrowed $37,000 from its bank
  3. 3) provided consulting services for $62,000 cash
  4. 4) paid back $27,000 of the bank loan
  5. 5) paid rent expense for $15,000
  6. 6) purchased equipment for $24,000 cash
  7. 7) paid $4,200 dividends to stockholders
  8. 8) paid employees' salaries of $33,000


What is Yowell's net income for Year 1?

Multiple Choice

  • $4,200

  • $27,000

  • $14,000

  • $59,800

5.Chow Company earned $5,500 of cash revenue, paid $3,000 for cash expenses, and paid a $1,200 cash dividend to its owners. Which of the following statements is true?

Multiple Choice

  • The net cash inflow from operating activities was $2,500.

  • The net cash outflow for investing activities was $1,200.

  • The net cash inflow from operating activities was $1,300.

  • The net cash outflow for investing activities was $2,500.

6. Packard Company engaged in the following transactions during Year 1, its first year of operations. (Assume all transactions are cash transactions.)

  1. 1) Acquired $1,150 cash from the issue of common stock.
  2. 2) Borrowed $620 from a bank.
  3. 3) Earned $800 of revenues cash.
  4. 4) Paid expenses of $290.
  5. 5) Paid a $90 dividend.

During Year 2, Packard engaged in the following transactions. (Assume all transactions are cash transactions.)

  1. 1) Issued an additional $525 of common stock.
  2. 2) Repaid $360 of its debt to the bank.
  3. 3) Earned revenues of $950 cash.
  4. 4) Incurred expenses of $440.
  5. 5) Paid dividends of $140.

Packard Company's net cash flow from financing activities for Year 2 is:

Multiple Choice

  • $360 outflow.

  • $500 inflow.

  • $25 inflow.

  • $385 outflow.
Answer #1

Part 2

The correct answer is

Inflow $9500

Cash flow from operations = consulting services cash revenue - paid rent expenses - employees salaries paid

= 44000 - 10500 - 24000

= $9500

Part 3

The correct answer is

$10000

Closing note payable balance = amount borrowed - amount repaid

= 40000 - 30000

= $10000

Part 4

The correct answer is

$14000

Net income = revenue - rent - Salaries

= 62000 - 15000 - 33000

= $14000

Part 5

The correct answer is

Net cash inflow from operating activities was $2500

Explanation

Cash inflow from operating activities

= cash revenue- cash expenses

= 5500 - 3000

= $2500

Part 6

The correct answer is

$ 25 inflow

Net cash inflow from financing activities for year 2

=Cash from issue of common stock- repaid debt by cash- paid dividend

=525 - 360 - 140

= $25